OpenStudy (anonymous):

2. Jackie, a 25 year old client, want to retire by age 65 with $1,500,000. How much would she have to invest annually, assuming a 6% rate of return?

OpenStudy (campbell_st):

this is an application of a geometric series.... the series is created from the annual investment being compounded a = 1.06 r = 1.06 and n = 40 P is the annual investment \[s _{40} = 1500000\] us the formula \[s _{n} =P \times (a \times(r^n -1))/(r -1)\] \[15000000 = P \times (10.6(1.06^40 -1))(1.06 -1)\] \[P = (1500000 \times 0.06)/(1.06^40 -1)\] the person needs to invest $ 9692.30

OpenStudy (campbell_st):

oops it should read \[1.06^{40}\]

OpenStudy (campbell_st):

where ever you see \[1.06^40\]